OpenStudy (mbussey):

He rowed 31/2 miles in an 1/2 hour. What is the average speed in miles per hour?

TheSmartOne (thesmartone):

remember the formula speed = distance/time so plug in the distance and time \(\Large Speed = \frac{\frac{31}{2}~miles}{\frac{1}{2}~hour}\) Multiply the numerator an denominator of the big fraction to get rid of the complex fraction and to get it in miles per hour So basically: \(\Large Speed = \frac{\frac{31}{2}~miles ~\times 2}{\frac{1}{2}~hour ~\times 2}= ?\)

OpenStudy (misssmartiez):

TheSmartOne is correct (medals). To make it easier, switch the fractions into decimals for easier calculation. 31 divided by 2 x 2 ----------- 1 divided by 2 x 2 . Example: 0.5 x 2 = 1 Now you find the answer for 31 divided by 2, and times THAT by 2. Got the two answers? Good. Now divide that, or leave it like it is. It should form a fraction;. Does the sheet ask for a fraction, or a whole number? - @MissSmartiez